# Using neuroevolution for designing soft medical devices

### v1

@@ 3.1 AFPO
This approach was conceived as an alternative method to avoid premature convergence in evolutionary algorithms [52], an issue exhibited by other evolutionary approaches. One of the fundamental aspects of AFPO is age, which, under this scope, can be understood as how long genotypic material has existed in the population [21]. Based on this, the age of a solution (i.e., an individual) is measured in generations.
The algorithm commences initialising the individuals composing the population at random, with age of all individuals set to one. If an individual survives to the next generation, its age is incremented by one. When genetic operators (i.e., crossover and mutation) take place, the age is inherited as the maximum age of the parents.
Since AFPO employs a single population, it tracks each individual’s fitness similarly to a standard evolutionary algorithm, including the genotypic age. The multi-objective optimisation procedure aims to determine the non-dominated Pareto front of the problem domain, where the objective is maximising fitness with minimum age. One way of visualising the mechanism of AFPO is the population (i.e., individuals) evolving in a two-dimensional plane of age and fitness. Figure 1 illustrates this dynamic.
@@ 3.2 NEAT
where $E$ is the number of excess genes, $D$ represents the disjoint genes, $\overline{W}$ is the mean weight differences of matching genes, $c_{1}$, $c_{2}$, $c_{3}$ are utilised to determine the relevance of $E$, $D$, and $\overline{W}$, respectively. Regarding $N$, it represents the number of genes in the individual with the largest number of genes; it is used to normalise the genome size during the evolutionary process. The distance $\delta$ determines in which species an individual is allocated utilising a compatibility threshold $\delta_{t}$. Individuals are allocated into the first species (taken at random) where their distance $\delta$ is $<\delta_{t}$, assuring that individuals do not belong to more than one species.
@@
The third issue was alleviated by minimising network structures, which is feasible due to the mechanism of speciation to protect innovation. Since NEAT can initialise individuals whose input neurons are fully connected to the output neurons (i.e., no hidden nodes are involved), new structures are incrementally generated when structures are mutated during the evolutionary process. The structures that exhibit a suitable performance can survive. The effect induced by this approach narrows down the search towards a minimal number of weight dimensions, which implies: (a) a reduction of runtime to find solutions [59] and (b) avoiding the generation and evaluation of complex structures during the evolutionary process.
@@ 3.3 HyperNEAT
    Activation functions. Typically, NEAT generates ANNs containing hidden nodes exclusively using the sigmoid function. HyperNEAT, nevertheless, allows the use of other activation functions, such as Gaussian, trigonometric, and periodic, in the nodes composing CPPNs. In this way, CPPNs can evolve, exploring a significantly more expansive search space of network topologies.
@@
    Substrate. HyperNEAT is capable of embodying the geometry of the domain problem due to the nature of CPPNs; therefore, the topology of ANNs is computed, taking into account their geometry. Thus, the geometric layout where HyperNEAT operates is called a substrate, which can be configured in diverse manners. One of the most popular configurations consists of a set of nodes allocated in a two-dimensional plane; it is known as grid. Another widely used configuration is the three-dimensional grid, which contains a set of nodes allocated in a three-dimensional space. Figure 3-a depicts an example of a grid, whereas Fig. 3-b presents an example of a three-dimensional grid.
@@
In general, the mechanism of HyperNEAT starts by defining the substrate. The next step consists of generating the initial population of CPPNs. Then, for each CPPN, the connections (i.e., weights) and bias of the substrate are calculated using Eqs. 2 and 3 for a two-dimensional grid, and Eqs. 4 and 5 for a three-dimensional grid. Once the substrate is built, its performance (and hence the performance of the CPPN) is evaluated, and a fitness score is assigned to the CPPN. After the evaluation stage, NEAT is applied to evolve the population of CPPNs. It is important to emphasise that the outputs of CPPNs are usually in the $[-1.0,1.0]$ range, which, in case it is necessary, implies a normalisation or mapping procedure.
Summarising, HyperNEAT aims to evolve the topology (i.e., the connectivity) and weights of ANNs that are represented by a substrate with a specific geometry through CPPNs.
4 Actuators generator
@@ 4.1 Experimental setup
The population is composed of 100 individuals (i.e., CPPNs) for the three approaches. Each evolutionary run lasted 3000 generations. The set of activation functions utilised for experimentation is the same as used in [28], and is composed of: sigmoid, sine, negative sine, square, negative square, square root of absolute, negative square root of absolute, absolute, and negative absolute. Activation functions are randomly chosen during CPPN initialisation and mutations. In order to perform a fair comparison for all approaches, the implemented initialisation procedure of the population is described in [28]. Moreover, for NEAT and HyperNEAT, the parameters employed for the evolutionary process of CPPNs are shown in Table 1. For each approach, 20 independent experimental runs were performed, by changing only the initial random populations.
@@
The fitness function consists of a physics engine called Voxelyze (also available in [2]), which simulates the physical response of morphologies under determined conditions. Morphologies are defined within Voxelyze by stacking voxels (the minimum building block in Voxelyze) of different materials together. Here two types of voxels were considered for simplicity of the designing landscape, namely one active voxel and one passive voxel. The original version of Voxelyze was modified (changes in the source code available in [1]) to trace the position of the catheter actuator in the $x$,$y$,$z$ axes during a time $t$. Furthermore, one end of actuators is fixed in such a way that they only move vertically (i.e., one degree of freedom) in the $yz$ plane. In addition, actuators have a passive enclosure to follow the findings of previous simulation works [60] and constraints of laboratory experiments, i.e. the bioreactor that a muscle actuator need to provide nutrients. The output of Voxelyze contains three values that describe the position of the actuator free tip (with respect to $x$,$y$,$z$ dimensions) for the initial position and the final one (after the predefined simulation time), and the number of voxels that compose the actuator.
@@
Since simulations require a significant amount of computation time, it is essential to try to reduce it. Thus, the implementation related to AFPO was written under the multiprocessing paradigm [28]. The implementation used for NEAT and HyperNEAT was designed under a client-server architecture to take advantage of its distributed computing capacities. The full description of the client-server implementation utilised in this research can be found in [4]. Finally, the hardware utilised for experimentation is: Processor: ARM (virtualised), nine cores (18 threads), 3.20GHz, RAM Memory: 16GB, LPDDR5.
@@ 4.1.1 NEAT configuration
Due to actuators being designed in a discrete three-dimensional layout, it is necessary to provide for each point across the layout: (i) the presence or not of a material voxel, and (ii) the material type the voxel is made. Thus, under NEAT, CPPNs are queried as follows:
@@
$Presence(x_{i},y_{i},z_{i})\begin{cases}\text{yes, }&|v_{i}|\geq 0.5\\ \text{no, }&\text{otherwise}\end{cases}$ (7)
@@
$Material~{}code(x_{i},y_{i},z_{i})\begin{cases}1\text{, }&|m_{i}|<0.5\\ 3\text{, }&\text{otherwise}\end{cases}$ (8)
@@ 4.1.2 HyperNEAT configuration
The first element of HyperNEAT to determine is the substrate. Thus, based on the discrete three-dimensional layout used to design actuators, the substrate has three neuron inputs. Furthermore, since the data required to determine the presence of a voxel in the $i$-$th$ point and the material of that voxel, the number of output neurons is two. Equation 9 presents how substrates are queried.
@@
where the tuple $(x_{i},y_{i},z_{i})$ is the coordinates of the $i$-$th$ point in the three-dimensional layout. The $v_{i}$ variable is associated with the presence or absence of a voxel in the $i$-$th$ point of the layout and is mapped using Eq. 7, and $m_{i}$ is related to the type of the voxel material in the point $i$ of the layout and is processed by Eq. 8. Figure 4 depicts the design of the two-dimensional substrate used for experimentation. The neuron allocation used for the substrate was found by conducting a series of experiments varying the number of neurons per hidden layer and the number of hidden layers in the $[1,5]$ range. Furthermore, the activation function implemented in the substrate is SELU due to its fast convergence and self-normalising properties [27].
@@
Once the substrate is defined, the following step defines how CPPNs are queried. Since the substrate was defined in a two-dimensional space, CPPNs are queried utilising Eq. 2 for calculating the weights among neurons, and Eq. 3 to compute the bias of the neurons allocated in the substrate. It is essential to point out that based on the definition of HyperNEAT [61], determining whether a connection is calculated is by a threshold. Thus, if $|weight_{mh}|>0.2$, it is normalised in the $[-3.0,3.0]$ range. Otherwise, there is no connection between the neuron $m$ and the neuron $h$.
@@ 4.2.1 Finding the fittest actuator
One vital feature of actuators is the displacement in the $yz$ plane, since they have one end fixed. This displacement can be understood as an indicator of the bending capacity exhibited by actuators: the more displacement the actuator reaches, the more bendable the catheter is. This capability is crucial because catheters can deliver drugs from diverse angles. Hence, the more bendable the actuator is, the higher the agility of the drug delivering catheter.
In this experiment, the evaluation of actuators is as follows: for each independent experimental run, 25 different controller phase offsets are generated at random and do not change over time. For each morphology of the population, the 25 controller phase offsets previously generated are used to simulate the displacement of the actuator. In other words, each morphology is simulated 25 times with different controller phase offsets. This procedure occurs each generation until the experimental run is complete. Thus, the aptitude ($apt$) of the actuator $a$ is calculated as follows:
@@
It is important to emphasise that this experiment does not consider the number of voxels as a penalty for the aptitude value.
@@ 4.2.2 Exploring the robustness of the fittest actuator
Each top actuator generated by all approaches is tested utilising 1000 different controller phase offsets of each active voxel. In other words, the controller phase offset scenarios were randomly generated and used by all the morphologies implied in this experiment (i.e., the three fittest actuator morphologies generated by each approach). Thus, the robustness of the actuator $a$ can be measured through a modified version of Eq. 10, which is defined as follows:
@@
When AFPO is applied (see Fig. 6-a), all morphologies exhibit significant differences, which is determined by testing all the data collected, and they are not normally distributed (Shapiro-Wilk test; $p<0.01$). Ten by the Kruskal-Wallis test, it is possible to confirm significant differences among the displacement observed in the three actuators ($p<0.01$). Thus, a performance ranking can be performed: Morphology 12 $>$ Morphology 19 $>$ Morphology 7 (Dunn’s test: $p<0.01$).
Under NEAT, all morphologies exhibit significant differences, and the data gathered are non-normal (Shapiro-Wilk, Kruskal-Wallis; $p<0.01$). Based on the previous result, it is feasible to produce a ranking of the performance of the three fittest actuators generated: Morphology 9 $>$ Morphology 16 $>$ Morphology 15 (Dunn’s test: $p<0.01$).
On the other hand, when HyperNEAT is applied, the three fittest morphologies generated show significant differences and non-normality is observed in the data gathered (Shapiro-Wilk, Kruskal-Wallis; $p<0.01$). Therefore, a ranking focused on performance can be generated: Morphology 2 $>$ Morphology 18 $>$ Morphology 11 (Dunn’s test: $p<0.01$).
@@
Furthermore, Fig. 7 shows violin plots comparing the performance of the fittest actuator morphology produced by AFPO, NEAT and HyperNEAT. Each violin presents the kernel density estimation of the frequency distribution, maximum, minimum and median values of 1000 controller phase offset scenarios. All the morphologies present significant differences, confirmed through the Kruskal-Wallis test ($p<0.01$). Based on this result, it is possible to rank the performance of the actuators as follows: NEAT $>$ HyperNEAT $>$ AFPO (Dunn’s test: $p<0.01$).
In general, results suggest that neuroevolution-based approaches, such as NEAT and HyperNEAT, are more adept at producing bendable actuators than those generated by AFPO, regardless of the controller phase offset scenario. This implies that NEAT and HyperNEAT have the potential to create more robust actuators than AFPO. Furthermore, despite the comparable performance of NEAT and HyperNEAT, the former has the edge in producing more robust and, consequently, more bendable actuators.
@@ 4.2.3 Identifying the minimum volume of the fittest actuators
Regarding completeness in the research presented, Fig. 8 depicts the fittest actuator morphologies generated by all approaches. The fittest actuator designed by AFPO (see Fig. 8-a) exhibits a pyramidal pattern consisting of having a few contractile voxels at the bottom and gradually incrementing the number of voxels towards the top of the actuator. Arguably, the lack of contractile support at the bottom of the actuator affects its bending properties. Furthermore, the fittest actuator generated by NEAT (see Figure 8-b) presents a striped diagonal pattern with no voxels throughout the contractile body composing the actuator. This pattern possibly allows the catheter actuator to bend suitably regardless of the controller phase offset. Finally, the fittest actuator created by HyperNEAT (see Figure 8-c) also exhibits a striped diagonal pattern with no voxels, only in the free end of the actuator. The striped pattern allows the actuator to bend; however, the solid contractile mass on the other end of the actuator arguably attenuates the bending movement.

### v2

@@ 3.1 AFPO
This approach was conceived as an alternative method to avoid premature convergence in some evolutionary algorithms [63]. One of the fundamental aspects of AFPO is age, which, under this scope, can be understood as how long genotypic material has existed in the population [25]. Based on this, the age of a solution (i.e., an individual) is measured in generations.
The algorithm commences initializing the individuals composing the population at random, with age of all individuals set to one. If an individual survives to the next generation, its age is incremented by one. When genetic operators (i.e., crossover and mutation) take place, the age is inherited as the maximum age of the parents.
Since AFPO employs a single population, it tracks each individual’s fitness similarly to a standard evolutionary algorithm, including the genotypic age. This multi-objective optimisation procedure aims to determine the non-dominated Pareto front of the problem domain, where the objective is maximizing fitness with minimum age. One way of visualizing the mechanism of AFPO is the population (i.e., individuals) evolving in a two-dimensional plane of age and fitness. Figure 1 illustrates this dynamic.
@@ 3.2 NEAT
where $E$ is the number of excess genes, $D$ represents the disjoint genes, $\overline{W}$ is the mean weight differences of matching genes, $c_{1}$, $c_{2}$, $c_{3}$ are utilised to determine the relevance of $E$, $D$, and $\overline{W}$, respectively. Regarding $N$, it represents the number of genes in the individual with the largest number of genes; it is used to normalize the genome size during the evolutionary process. The distance $\delta$ determines in which species an individual is allocated utilising a compatibility threshold $\delta_{t}$. Individuals are allocated into the first species (taken at random) where their distance $\delta$ is $<\delta_{t}$, assuring that individuals do not belong to more than one species.
@@
The third issue was alleviated by minimizing network structures, which is feasible due to the mechanism of speciation to protect innovation. Since NEAT can initialize individuals whose input neurons are fully connected to the output neurons (i.e., no hidden nodes are involved), new structures are incrementally generated when structures are mutated during the evolutionary process. The structures that exhibit a suitable performance can survive. The effect induced by this approach narrows down the search towards a minimal number of weight dimensions, which implies: (a) a reduction of runtime to find solutions [73] and (b) avoiding the generation and evaluation of complex structures during the evolutionary process.
@@ 3.3 HyperNEAT
    Activation functions. Typically, NEAT generates ANNs containing hidden nodes exclusively using the sigmoid function. HyperNEAT, nevertheless, allows the use of other activation functions, such as Gaussian, trigonometric, and periodic, in the nodes composing CPPNs. In this way, CPPNs can evolve, exploring a significantly more expansive search space of network functionality.
@@
    Substrate. HyperNEAT is capable of embodying the geometry of the domain problem due to the nature of CPPNs; therefore, the topology of ANNs is computed, taking into account their geometry. Thus, the geometric layout where HyperNEAT operates is called a substrate, which can be configured in diverse manners. One of the most popular configurations consists of a set of nodes allocated in a two-dimensional plane, known as grid. Another widely used configuration is the three-dimensional grid, which contains a set of nodes allocated in a three-dimensional space. Figure 3-a depicts an example of a grid, whereas Fig. 3-b presents an example of a three-dimensional grid.
@@
In general, the mechanism of HyperNEAT starts by defining the substrate. The next step consists of generating the initial population of CPPNs. Then, for each CPPN, the connections (i.e., weights) and bias of the substrate are calculated using Eqs. 2 and 3 for a two-dimensional grid, and Eqs. 4 and 5 for a three-dimensional grid. Once the substrate is built, its performance (and hence the performance of the CPPN) is evaluated, and a fitness score is assigned to the CPPN. After the evaluation stage, NEAT is applied to evolve the population of CPPNs. It is important to emphasize that the outputs of CPPNs are usually in the $[-1.0,1.0]$ range, which, in case it is necessary, implies a normalization or mapping procedure.
Summarizing, HyperNEAT aims to evolve the topology (i.e., the connectivity) and weights of ANNs that are represented by a substrate with a specific geometry through CPPNs.
4 Experimental setup and Results
@@ 4.1 Experimental setup
The population is composed of 100 individuals (i.e., CPPNs) for the three approaches. Each evolutionary run lasted 3000 generations. The set of activation functions utilised for experimentation is the same as used in [34], and is composed of: sigmoid, sine, negative sine, square, negative square, square root of absolute, negative square root of absolute, absolute, and negative absolute. Activation functions are randomly chosen during CPPN initialization and mutations. In order to perform a fair comparison for all approaches, the implemented initialization procedure of the population is described in [34]. Moreover, for NEAT and HyperNEAT, the parameters employed for the evolutionary process of CPPNs are shown in Table 1. For each approach, 20 independent experimental runs were performed, by changing only the initial random populations.
@@
Since simulations require a significant amount of computational time, it is essential to mitigate this limitation. Thus, the implementation related to AFPO was previously written under the multiprocessing paradigm [34] and was used here under the same principle. The implementation for NEAT and HyperNEAT was designed under a client-server architecture to take advantage of distributed computing capacities [4]. The hardware utilised for experimentation in this study is described as follows: Processor: ARM (virtualised), nine cores (18 threads), 3.20GHz, RAM Memory: 16GB, LPDDR5. Provided that the evolutionary optimisation is implemented under a client-server architecture, the utilisation of High-performance computing (HPC) can significantly accelerate this methodology, so that it complies with time constraints of real-world optimisation problems. Additionally, some alternations in the algorithmic implementation, like the initialization methodology, have proved to contribute with an acceleration of 30% on the computations [73].
@@ 4.1.1 NEAT configuration
Due to actuators being designed in a discrete three-dimensional layout, it is necessary to provide for each point across the layout: (i) the presence or not of a material voxel, and (ii) the material type of each voxel. Thus, under NEAT, CPPNs are queried as follows:
@@
$Presence(x_{i},y_{i},z_{i})=\begin{cases}\text{yes, }&|v_{i}|\geq 0.5\\ \text{no, }&\text{otherwise}\end{cases}$ (7)
@@
$Material~{}code(x_{i},y_{i},z_{i})=\begin{cases}1\text{, }&|m_{i}|<0.5\\ 3\text{, }&\text{otherwise}\end{cases}$ (8)
@@ 4.1.2 HyperNEAT configuration
The first element of HyperNEAT to determine is the substrate. Thus, based on the discrete three-dimensional layout used to design actuators, the substrate has three neuron inputs. Furthermore, since the data required to determine are the presence of a voxel in the $i$-$th$ point and the material of that voxel, the amount of output neurons is two. Equation 9 presents how substrates are queried.
@@
where the tuple $(x_{i},y_{i},z_{i})$ is the coordinates of the $i$-$th$ point in the three-dimensional layout. The $v_{i}$ variable is associated with the presence or absence of a voxel in the $i$-$th$ point of the layout and is mapped using Eq. 7, and $m_{i}$ is related to the type of the voxel material in the point $i$ of the layout and is processed by Eq. 8. Figure 4 depicts the design of the two-dimensional substrate used for experimentation. The neuron allocation used for the substrate was decided by conducting a series of experiments varying the number of neurons per hidden layer and the number of hidden layers in the $[1,5]$ range. Furthermore, the activation function implemented in the substrate is SELU due to its fast convergence and self-normalizing properties [33].
@@
Once the substrate is defined, the following step defines how CPPNs are queried. Since the substrate was defined in a two-dimensional space, CPPNs are queried utilising Eq. 2 for calculating the weights among neurons, and Eq. 3 to compute the bias of the neurons allocated in the substrate. It is essential to point out that based on the definition of HyperNEAT [75], determining whether a connection is calculated is by a threshold. Thus, if $|weight_{mh}|>0.2$, it is normalized in the $[-3.0,3.0]$ range. Otherwise, there is no connection between the neuron $m$ and the neuron $h$.
@@ 4.2.1 Finding the fittest actuator
One vital feature of actuators is the displacement in the $yz$ plane, since they have one end fixed. This displacement can be understood as an indicator of the bending capacity exhibited by actuators: the more displacement the actuator reaches, the more bendable the catheter is.
For instance, if a soft actuator exhibits no displacement despite a determined stimulation, the angular range where it can operate is $[0,\alpha]$, where $\alpha=0^{\circ}$, which can be interpreted as the soft actuator presents a poor performance. On the other hand, if the movement observed in the soft actuator allows to satisfy $\alpha>0^{\circ}$, then the soft actuator can operate in a more comprehensive angular range (i.e., a better performance). In general, the bigger the displacement of the soft actuator (and hence $\alpha$) is, the more comprehensive the angular range where the soft actuator can operate.
In this experiment, the evaluation of actuators is as follows: for each independent experimental run, 25 different controller strategies (or phase offsets in the oscillations of the simulated active voxels) are generated at random and do not change throughout the evolution. For each morphology of the population, the 25 controller phase offsets previously generated are used to stimulate the displacement of the actuator. In other words, each morphology is stimulated 25 times with different controller phase offsets. This procedure occurs for each generation until the experimental run is complete. Thus, the aptitude $apt$ of the actuator $a$ is calculated as follows:
@@
It is important to emphasize that this experiment does not consider the number of voxels as a penalty for the aptitude value.
@@ 4.2.2 Exploring the robustness of the fittest actuator
Each top actuator generated by all approaches is tested utilising 1000 different controller phase offsets of each active voxel. In other words, the controller phase offset scenarios were randomly generated and used by all the morphologies studied in this experiment (i.e., the three fittest actuator morphologies generated by each approach). Thus, the robustness (i.e., the aptitude $apt$) of the actuator $a$ can be measured as follows:
@@
When AFPO morphologies are considered (see Fig. 6-a), all morphologies exhibit significant differences, a fact that is determined by testing all the data collected, and realizing that they are not normally distributed (Shapiro-Wilk test; $p<0.01$). Then by the Kruskal-Wallis test, it is possible to confirm significant differences among the displacement observed in the three actuators ($p<0.01$). Thus, a performance ranking can be performed: Morphology 12 $>$ Morphology 19 $>$ Morphology 7 (Dunn’s test: $p<0.01$).
Under NEAT, all morphologies exhibit significant differences, and the data gathered are non-normal (Shapiro-Wilk, Kruskal-Wallis; $p<0.01$). Based on the previous result, it is feasible to produce a ranking of the performance of the three fittest actuators generated by NEAT: Morphology 9 $>$ Morphology 16 $>$ Morphology 15 (Dunn’s test: $p<0.01$).
On the other hand, when HyperNEAT results are analysed, the three fittest morphologies generated show significant differences and non-normality is observed in the data gathered (Shapiro-Wilk, Kruskal-Wallis; $p<0.01$). Therefore, a ranking focused on performance can be assumed: Morphology 2 $>$ Morphology 18 $>$ Morphology 11 (Dunn’s test: $p<0.01$).
@@
To conclude, Fig. 7 shows violin plots comparing the performance of the fittest actuator morphology produced by AFPO, NEAT and HyperNEAT. Each violin presents the kernel density estimation of the frequency distribution, maximum, minimum and median values of the 1000 controller phase offset scenarios. All the morphologies present significant differences, confirmed through the Kruskal-Wallis test ($p<0.01$). Based on this result, it is possible to rank the performance of the actuators as follows: NEAT $>$ HyperNEAT $>$ AFPO (Dunn’s test: $p<0.01$).
In general, results suggest that neuroevolution-based approaches, such as NEAT and HyperNEAT, are more adept at producing bendable actuators than those generated by AFPO, regardless of the controller phase offset scenario. This realization comes as a contradiction to the previous test, where AFPO seemed as the most appropriate method, whereas, only 25 phase offsets were tested during the optimisation evolutionary process. This implies that NEAT and HyperNEAT have the potential to create more robust actuators than AFPO. Furthermore, despite the comparable performance of NEAT and HyperNEAT, the former has the edge in producing more robust and, consequently, more bendable actuators. As mentioned previously, HyperNEAT can prove advantageous over problems that have a strong geometric regularity, based on its ability to evolve network structures over geometric substrates. On the other hand, the NEAT algorithm provided better results in this robustness analysis. This can be attributed to two factors. Firstly, the HyperNEAT substrate was only fine-tuned in a limited range of (i) hidden layers and (ii) neurons per hidden layer (i.e., [1,5] for both), because of the high computational requirement related with testing higher ranges. As a result, the best substrate configuration for the given problem can be more complicated than the ones tested within this study. Secondly, NEAT is based on evolving CPPNs (CPPN-NEAT), that have been proved to be very efficient in emerging behaviors that can facilitate movement [14].
@@ 4.2.3 Identifying the minimum volume of the fittest actuators
For illustration reasons, Fig. 8 depicts the fittest actuator morphologies generated by all three approaches. The fittest actuator designed by AFPO (see Fig. 8-a) exhibits a pyramidal pattern consisting of having a few contractile voxels at the bottom and gradually incrementing the number of voxels towards the top of the actuator. Arguably, the lack of contractile support at the bottom of the actuator affects its bending properties. Furthermore, the fittest actuator generated by NEAT (see Figure 8-b) presents a striped diagonal pattern with no voxels throughout the contractile body composing the actuator. This pattern possibly allows the catheter actuator to bend suitably regardless of the controller phase offset. Finally, the fittest actuator created by HyperNEAT (see Figure 8-c) also exhibits a striped diagonal pattern with no voxels, only in the free end of the actuator. The striped pattern allows the actuator to bend; however, the solid contractile mass on the other end of the actuator arguably attenuates the bending movement.
To sum up, the NE-based methodologies have produced the most robust actuator morphologies, that directly relates to a high chance of performing efficiently on real-world applications. The random phase offsets used in this study can be regarded to assess morphologies for noise resistance in the simulations that is a way to trivially predict the success of their physical counterpart when manufactured in a real-world application [28]. The superior performance of AFPO on finding the fittest morphology can be preferred only in cases that the actuators can be controlled with high accuracy, due to the small variety of phase offsets tested throughout the evolutionary optimisation (for computational efficiency reasons). The proposed methodology can be instantiated with the parameters of the building blocks (or physical characteristics of available materials) to be utilised for the soft actuator manufacturing process, the physical constraints of the problem, the physical description of the environment that will accommodate the actuator and a target behavior (or fitness function). Then, through the evolutionary optimisation of the NE-based algorithms, an efficient morphology of the actuator will be produced, that will be in the form of what type of material should be placed at each point in the three-dimensional design space.
4.3 Error analysis
To understand the variability of the performance of actuators and, consequently, of the algorithms used in this research, two experiments are performed: (i) varying the properties of two simulated materials and (ii) altering the control strategy used in previous experiments. It is important to emphasize that the fittest actuator morphology found in the experiments described in Section 4.2.2 was utilised for the experiments described in the following sections. Namely, Morphology 9, found by NEAT (N-M9).
4.3.1 Varying material properties
This experiment changes the value of two mechanical properties of the contractile material used for the actuator morphologies during simulations. These two properties are:
  1. 1.
    Young’s modulus. It describes the stiffness of materials. This property is related to the elasticity of materials under compression or tension [29].
  2. 2.
    Poisson’s ratio. It is the negative ratio of transverse strain to axial strain. In other words, this property is the deformation of a material perpendicular to the load direction [29].
During the previous experiments (see Section 4), Young’s modulus and Poisson’s ratio were set to 5e6 and 0.35, respectively. In this experiment, each value is varied in the following range: $[-10\%,-5\%,5\%,10\%]$.
Figure 9 depicts violin plots comparing the displacement exhibited (i.e., the performance) by N-M9 varying the values of the Young’s modulus and the Poisson’s ratio. Each violin plot shows the median, minimum, maximum and kernel density estimation of the frequency distribution under 1000 different offset controller scenarios.
Figure 9: Performance of N-M9 varying the values of: (a) Young’s modulus, and (b) Poisson’s ratio.
When the value of Young’s modulus is altered (see Figure 9-a), significant differences can be observed among the performances. All the data gathered are non-normally distributed (Shapiro-Wilk; $p<0.01$). It is possible to confirm significant differences between the performances whose Young’s value is $<{5e6}$ (Dunn’s test; $p<0.01$). Furthermore, significant differences are exhibited when these two performances are compared against those whose Young’s value is $\geq{5e6}$ (Dunn’s test; $p<0.01$). However, there are no significant differences among the performances whose Young’s value is $\geq{5e6}$ (Dunn’s test; $p>0.05$).
As expected, the results indicate that the material’s stiffness significantly influences the actuator morphology’s performance. In general, the less stiff the material is, the less controllable it is, and hence, the less performance is observed. On the other hand, if the material is stiffer, there is no significant improvement in performance.
Furthermore, when the value of the Poisson’s ratio is varied, there are no significant differences among the performances. This is confirmed by first testing the collected data using the Shapiro-Wilk test and obtaining $p<0.01$, which indicates that the data is non-normally distributed. Then, applying Dunn’s test, the result is $p>0.05$, suggesting that the deformation of the material does not significantly affect the performance of the actuator morphology.
4.3.2 Testing the control strategy
In this experiment, a random controller phase offset is used. This controller phase offset is copied 18 times. Each copy represents a scenario where a slice in the plane $yz$ of N-M9 is affected. The controller phase offset values associated with the voxels composing the slice are set to 0. It is essential to point out that although the dimension of N-M9 in the $x$ axis is 20, only 18 slices are available due to the passive enclosure embodied (see Section 3). For visual reference, see Figure 8-b. Furthermore, the value of Young’s modulus and Poisson’s ratio are set to 5e6 and 0.35 respectively (see Section 4.3.1).
Figure 10: Performance of N-M9 under 18 different scenarios where the values of the controller phase offset associated with the $n$ slice are set to 0.
Figure 10 presents a bar plot showing the displacement reached by N-M9 when the values of the controller phase offset associated with $n$ slice are set to 0 (blue bars). In addition, a green bar shows the performance of N-M9 with no modification in the values of the controller phase offset; it is used as a baseline.
As expected, no significant variation in terms of performance is observed in the slices that are near the fixed end of N-M9. This behavior occurs from slices 1 to 14. Furthermore, those slices that are near the free end of N-M9 exhibit a significant variation, which is observed in slices 16 to 18.
In general, despite the specific noise applied to the controller input, the performance of the actuator morphology is consistent (i.e., there are no noticeable variations) in the vast majority of scenarios. On the other hand, the variations in performance observed are arguably due to external interactions that affect the material’s mechanical properties, such as the absence of physical support (i.e., no fixed end).